OpenStudy (anonymous):

y = 7/x^2 + 10 Find f(x) and g(x) so the function can be expressed as y = f(g(x)).

OpenStudy (anonymous):

help me!

OpenStudy (anonymous):

at first care that these functions are not uniqe... one pair of them is:

OpenStudy (anonymous):

f(x) = 7/x + 10 and g(x) = x^2 need explanation?

OpenStudy (anonymous):

if f(x) =7/x+10 and g(x)=x^2 f(g(x))=7/x^2+10